# Total cross section of the process $e^{+}+e^{-}\to\Sigma^{+}+\bar{\Sigma}^{-}$ including the $D$-meson loop and three-gluon contributions

### v1

@@ Abstract
In the present paper, we investigate the production of a baryon pair in $e^{+}e^{-}$ annihilation to study the structure of baryons. To study the basic structure of the Standard Model, it is also necessary to consider the baryon-antibaryon pair production at the electron-positron linear collider. The production of a baryon pair in electron-positron annihilation provides a powerful tool at higher center-of-mass energies. In this work, we present phenomenological results for $\Sigma^{+}\bar{\Sigma}^{-}$ production in $e^{+}e^{-}$ collision at the BESIII Collider. We investigate a hyperon pair produced in the reaction $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$. Two different contributions are considered: the $D$ meson loop and the three-gluon charmonium annihilation loop. Taking into account the contributions of the $D$-meson loop and three-gluon loops and the interference of all diagrams, we calculate the total cross section of the process $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$. With respect to a purely electromagnetic mechanism, large relative phases are generated for these contributions. We obtain as a byproduct a fit of the electromagnetic form factor of the $\Sigma$ hyperon for a large momentum transferred region. In this process, in addition to $\psi(3770)$ charmonium, we also take into account the contributions of other charmonium (-like) states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$. The obtained results are in good agreement with experimental data.
@@ I Introduction
It is understandable that the nucleons, as the lightest baryons, are the largest component of the observable matter in the Universe and are shown to be non-pointlike particles A1 ; A2 . In fact, the hyperons are the SU(3)-flavor-octet partners of the nucleons that contain one or more strange quarks and necessary crucial additional dimensions to study nucleon structures A3 ; A4 . It should be noted that to understand the strong interaction where quarks form hadrons is one of the most challenging questions in contemporary physics. It is known that electromagnetic form factors (EMFFs) are fundamental observables of baryons that are closely related to their internal structure and dynamics. That is, EMFFs, which parameterize the internal structure and dynamics of hadrons, are fundamental observables of hyperons and nucleons for understanding the effects of Quantum Chromodynamics (QCD) in hadron resonances A5 ; A6 ; A7 ; A8 .
It must be noted that electromagnetic form factors describe well the changes at the point photon-hadron vertex due to the internal structure and dynamics of hadrons. Therefore, our knowledge of the electromagnetic form factors of the nucleon has significantly improved over the last years. We can say that the electromagnetic form factors, can be investigated using the process of electron-positron interaction. Really, it is possible to obtain valuable information about their structure from the electromagnetic form factors of nucleons with high momentum transfer. We can say that the internal structure of baryons is parameterized using electromagnetic form factors.
@@
Like others, the vector ($J^{PC}=1^{--}$) charmonium-like $Y$-family states, e.g., $Y(4360)$, were firstly observed and confirmed in the process $e^{+}e^{-}\to(\gamma_{ISR})\pi^{+}\pi^{-}\psi(3686)$ by BaBar and Belle experiments A12 ; A13 . However, the nature of $Y(4360)$ has remained mysterious.
The $Y(4660)$ has been reported in ISR production in the process $e^{+}e^{-}\to(\gamma_{ISR})\psi(2S)\pi^{+}\pi^{-}$ A14 ; A15 . It was known that $Y(4660)$ was observed only in the Belle experiment A15 and, therefore, it is important to confirm the existence of this state.
@@
It is known that the study of the physics of charm plays an important role in understanding the standard model. It should be noted that experiments with electron-positron colliding beams have provided a number of interesting data on the annihilation cross section in the vector meson region, as well as some preliminary data in the region of several GeV.
It is necessary to note that for studying the strong interaction in the transition region between the non-perturbative and perturbative regimes of QCD hyperons are ideal probes. Also, in studying the symmetry properties in particle physics, weak decays of two-particle hyperons play an important role.
The region in the center of mass with energies above the threshold for the production of the charmonium is of great interest for the theory due to its richness of $c\bar{c}$ states, the properties of which have not been sufficiently studied. The study of mesons and baryons containing a few one-charm quarks is called open charm physics.
We know that a lot of charmonia have been observed since 1974 A18 ; A19 , but this is not the end of the whole history of the charmonium family, especially of observations of charmonium-like XYZ states A20 ; A21 .
It is clear that the charmonium family group, by adding more states, may provide important information for understanding the nonperturbative behavior of the strong interaction, which is intensively associated with color restriction in quantum chromodynamics (QCD). The vector charmonium-like $Y$ state plays a special role in the construction of the charmonium family.
@@
We also would like to note that in experimental measurements, all charmonium states were observed below the threshold of a charmed pair of quarks in mass ($2m_{D}$), and it was shown that their observed spectrum is consistent with the prediction of the model of the potential of charm and anti-charm A27 .
Two-particle baryon decays of vector ($J^{PC}=1^{--}$) charmonium (-like) resonances provide a rich laboratory for predictions of quantum chromodynamics A28 ; A29 .
@@
For the process $e^{+}e^{-}\to\bar{p}p$, the authors of the work Bianconi noted that there are deviations in the existence in the BABAR data in time-like proton form factors from the pointlike behavior of proton-antiproton electromagnetic current. Namely, in this paper $F_{0}$ + $F_{osc}$ was used in terms of the form factor where $F_{0}$ is expressed as the long-term trend of the form factor, and $F_{osc}$ is a function that has the form exp(-Bp)cos(Cp) where $p$ is the relative momentum of the $\bar{p}p$ pair in the final state.
@@
In these experimental works I5 ; I6 ; I7 ; I8 ; I9 ; I10 ; I11 ; I12 ; I13 , from the point of view of available data, there is an electromagnetic form factor in a time-like region, which is consistent with naive rules for counting quarks and the prediction of perturbative QCD (pQCD) at large $q^{2}$ I14 ; I15 . Therefore, in these cases, to provide a unique opportunity to study the internal structure of hadrons and the electromagnetic properties of hyperons, timelike form factors are used.
It should be noted that the data listed in I16 , which are obtained mainly from experiments on electron-positron and proton-proton (proton-antiproton) interactions, determine today most of our knowledge about experimentally established nucleon resonances.
The charmonium resonance mass is present in the transition region between the perturbative and nonperturbative regimes of the $\psi(3770)$. The investigation of the production of a baryon pair in the $e^{+}e^{-}$ annihilation allows us to study the structures of baryons. To gain knowledge about the structure in perturbative and nonperturbative strong interactions, it is necessary to study the strong and hadronic decays of $\psi(3770)$ in this energy region I17 .
After analyzing these samples together with the data, it became possible to conduct a theoretical study taking into account the interference of resonant and non-resonant amplitudes for the exclusive decay of $\psi(3770)$. It is known that for several years BESIII has collected the largest sample of data on e+e collisions at an energy of 3.773 GeV.
@@
What follows is the decay process in these charmonium states, i.e., the decay into light hadrons via either a three-gluon process ($e^{+}e^{-}\to\psi\to ggg\to hadrons$) or a one-photon process ($e^{+}e^{-}\to\psi\to\gamma^{*}\to hadrons$). This is followed by a decay in these charmonium states, i.e., decay into light hadrons through either the three-gluon process ($e^{+}e^{-}\to\psi\to ggg\to hadrons$), or the one-photon process ($e^{+}e^{-}\to\psi\to\gamma^{*}\to hadrons$).
Really, according to the Okubo-Zweig-Iizuki rule (OZI), it is expected that $\psi(3770)$, the lowest charmonium state lying at $1^{--}$ above the threshold of $D\bar{D}$, will predominantly decay into the $D\bar{D}$ final states I18 ; I19 . Binary systems with two particles, which are obtained in finite states, make it possible to further simplify the consideration of processes with charmonium in the intermediate state I20 .
To test a QCD prediction that can be understood in terms of the amplitudes of the quark distribution in hadron-hadron pairs and the conservation of the total hadron helicity, it is necessary to study the process of $\psi(3770)$ production in e+e annihilation and its subsequent decay into two hadrons.
@@
It can be expected that the resonance of $\psi(3770)$ will almost completely decay into pure $D\bar{D}$ I21 , and in the baryon sector at the electron-positron collider, the production of a baryon-anti-baryon pair can be tested by fundamental symmetries, in particular, when the probability of the process increases due to the resonance such as the $J/\psi$ (or $\psi$)I22 .
It should be noted that the BESIII collaboration has until now conducted highly-accurate studies of the possible threshold increase in the $e^{+}e^{-}\to\Sigma^{\pm}\bar{\Sigma}^{\mp}$ I23 , $\Xi^{-}\bar{\Xi}^{+}$ I24 ; I24A and $\Sigma^{0}\bar{\Sigma}^{0}$ I25 processes and has also shown that the cross section does not vanish near the threshold.
It can be understood that the threshold effect thus obtained will be useful for measuring the near-threshold pair generation of hyperons in the production of $\Sigma^{+}\bar{\Sigma}^{-}$ in the $e+e-$ annihilation process, which was observed in the BESIII experiment I26 .
@@
In this work, our first purpose is to calculate the total cross section in the framework of QED and Standard Model (SM) taking into account the $D$-meson loop and three-gluon contributions.
Our aim in this paper is to study the characteristics of the $\Sigma^{0}\bar{\Sigma}^{0}$ production in the $e^{+}e^{-}$ annihilation process, in except for the one-photon exchange the exchange of several charmonium(-like) states, $\psi(3770),\,\psi(4040),\,\psi(4160),\,Y(4230),\,Y(4360),\,\psi(4415),\,\psi(4660)$, which decay into $\Sigma^{+}\bar{\Sigma}^{-}$ in the final state with taking into account the contribution of the $D$-meson loop and three-gluon loops.
@@ II The process $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ in Born approximation
$d\Phi_{2}=(2\pi)^{4}\delta(p_{+}+p_{-}-q_{+}-q_{+})\frac{d\bf{q_{+}}}{(2\pi)^{3}2E_{+}}\frac{d\bf{q_{+}}}{(2\pi)^{3}2E_{+}}=$
$=\frac{1}{(2\pi)^{2}}\delta(p_{+}+p_{-}-q_{+}-q_{-})\frac{d\bf{p_{+}}}{2E_{+}}\frac{d\bf{p_{-}}}{2E_{-}}=\frac{|\bf{p}|}{16\pi^{2}\sqrt{s}}\,d\Omega_{\Sigma}=\frac{\beta}{16\pi}\,d\cos\theta_{\Sigma},$ (5)
where $d\Omega_{\Sigma}=d\phi_{\Sigma}\,d\cos\theta_{\Sigma}=2\pi\,d\cos\theta_{\Sigma}$, and $\phi_{\Sigma}$ and $\theta_{\Sigma}$ are the azimuthal and polar angles of the $\Sigma$-hyperon momentum in the $e^{+}e^{-}$ rest frame, that is $\theta_{\Sigma}$ is the angle between the directions of the momenta of the electron $\bf{p_{2}}$ and $\Sigma$-hyperon $\bf{q}_{1}$ (Fig. 1). We note that the modulus of three momenta of the final $\Sigma^{+}$-hyperon (or $\bar{\Sigma}^{-}$-hyperon) is defined in this case by the $\delta$-function in the phase volume, i.e.,
$|{\bf{p}}|\equiv|{\bf{q_{1}}}|=|{\bf{q_{2}}}|=\frac{\sqrt{s}}{2}\sqrt{1-\frac{4M^{2}_{\Sigma}}{s}}=\frac{\sqrt{s}}{2}\beta,$ (6)
@@
where $C$ is a free parameter that fits the experimental data and $\Lambda_{QCD}$ is the QCD scale parameter. We want to note that the constant $C$ should be fitted for the hyperon-anti-hyperon production according to specific experimental data in the energy range of the corresponding experiment.
In the present work for the $\Sigma^{+}\bar{\Sigma}^{-}$ pair production in the $e^{+}e^{-}$ annihilation process (1), we fix this constant $C$ using the BESIII measurement I26 , and in Fig. 3 we show the dependence of the total cross section as a function of the center-of-mass energy $\sqrt{s}$. When the value is $\Lambda$ = 300 MeV in the Born cross section from (11) for the parameter $C$, after fitting with respect to these data we obtain the following value:
@@ III The quarkonium $\psi(3770)$ intermediate state
but only with a different constant $g_{e}=F_{1}^{\psi\to\Sigma^{0}\bar{\Sigma}^{0}}(M_{\psi}^{2})$, that depends on the charmonium mass and is equal to the value of the form factor $F_{1}^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}(M_{\psi}^{2})$ which is the value of the form factor for the vertex $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ on the $\psi(3770)$ mass shell [we note that here we accept the same approximation as in the Born case and assume that $F_{2}^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}(M_{\psi}^{2})=0]$. We know from I16 the total decay width of $\psi\to e^{+}e^{-}$, which is equal to $\Gamma_{\psi\to e^{+}e^{-}}=261\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{eV}$, and we can calculate this constant $g_{e}$,
@@
It is shown Kuraev that the imaginary part of the vertex $e\bar{e}\to\psi$ is small and it is less than 10 % of the real part. In this case, we neglect such a possible imaginary part.
@@
$\approx-eg_{e}\,\mbox{Sp}[\hat{p_{-}}\gamma_{\mu}\hat{p_{+}}\gamma^{\mu}]\approx 4\,e\,g_{e}s.$ (23)
@@
In this formula (24), the subintegral expression contains the dynamics of the transformation of the charmonium into the $\Sigma^{0}\bar{\Sigma}^{0}$ pair, and therefore, it can be written in a separate form
$S_{i}(s)=\frac{eg_{e}\beta}{48\pi s}\int\limits_{-1}^{1}d\cos\theta_{\Sigma^{0}}\sum_{s^{\prime}}(J_{\gamma\to\Sigma^{0}\bar{\Sigma}^{0}}^{\alpha})^{*}J^{\psi\to\Sigma^{0}\bar{\Sigma}^{0}}_{\alpha}.$ (25)
Thus, if we take into account formula (25) formula in the (24) for the total cross section, then the expression for the interference contribution can be written in the following form:
@@
In order to reconstruct the total cross section, we use the procedure described in Bystritskiy [equations (15) and (16)] and with the help of the interference contribution (26) with the total relative phase between the Born amplitude $\mathcal{M}_{B}$ and the amplitude with the charmonium contribution $\mathcal{M}_{\psi}$.

### v3

@@ Abstract
In the present paper, we investigate the production of a baryon pair in $e^{+}e^{-}$ annihilation to study the structure of baryons. To study the basic structure of the Standard Model, it is also necessary to consider the baryon-antibaryon pair production at the electron-positron linear collider. The production of a baryon pair in electron-positron annihilation provides a powerful tool at higher center-of-mass energies. In this work, we present phenomenological results for $\Sigma^{+}\bar{\Sigma}^{-}$ production in $e^{+}e^{-}$ collision at the BESIII Collider. We investigate a hyperon pair produced in the reaction $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$. Two different contributions are considered: the $D$ meson loop and the three-gluon charmonium annihilation loop. Taking into account the contributions of the $D$-meson loop and three-gluon loops and the interference of all diagrams, we calculate the total cross section of the process $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$. With respect to a purely electromagnetic mechanism, large relative phases are generated for these contributions. We obtain as a byproduct a fit of the electromagnetic form factor of the $\Sigma$ hyperon for a large momentum transferred region. In this process, in addition to $\psi(3770)$ charmonium, we also take into account the contributions of other charmonium (like) states, $\psi(4040)$,   $\psi(4160)$,   $Y(4230)$,  $Y(4360)$,  $\psi(4415)$, and $Y(4660)$. The obtained results are in good agreement with experimental data.
@@ I Introduction
It is understandable that the nucleons, as the lightest baryons, are the largest component of the observable matter in the Universe and are shown to be nonpointlike particles A1 ; A2 . In fact, the hyperons are the SU(3)-flavor-octet partners of the nucleons that contain one or more strange quarks and necessary crucial additional dimensions to study nucleon structures A3 ; A4 ; A41 ; A42 . It should be noted that to understand the strong interaction where quarks form hadrons is one of the most challenging questions in contemporary physics. It is known that electromagnetic form factors (EMFFs) are fundamental observables of baryons that are closely related to their internal structure and dynamics. That is, EMFFs, which parameterize the internal structure and dynamics of hadrons, are fundamental observables of hyperons and nucleons for understanding the effects of Quantum Chromodynamics (QCD) in hadron resonances A5 ; A6 ; A7 ; A8 .
It must be noted that electromagnetic form factors describe well the changes at the point photon-hadron vertex due to the internal structure and dynamics of hadrons. Therefore, our knowledge of the electromagnetic form factors of the nucleon has significantly improved over the last years. We can say that the electromagnetic form factors, can be investigated using the process of electron-positron interaction. Really, it is possible to obtain valuable information about their structure from the electromagnetic form factors of nucleons with high momentum transfer.
@@
Like others, the vector ($J^{PC}=1^{--}$) charmoniumlike $Y$-family states, e.g., $Y(4360)$, were firstly observed and confirmed in the process $e^{+}e^{-}\to(\gamma_{ISR})\pi^{+}\pi^{-}\psi(3686)$ by $BABAR$ and Belle experiments A12 ; A13 . However, the nature of $Y(4360)$ has remained mysterious.
The $Y(4660)$ has been reported in ISR production in the process $e^{+}e^{-}\to(\gamma_{ISR})\psi(2S)\pi^{+}\pi^{-}$ A121 ; A131 . It was known that $Y(4660)$ was observed only in the Belle experiment A131 and, therefore, it is important to confirm the existence of this state.
@@
It is known that the study of the physics of charm plays an important role in understanding the standard model. It should be noted that experiments with electron-positron colliding beams have provided a number of interesting data on the annihilation cross section in the vector meson region, as well as some preliminary data in the region of several GeV. It is necessary to note that for studying the strong interaction in the transition region between the nonperturbative and perturbative regimes of QCD hyperons are ideal probes.
The region in the center of mass with energies above the threshold for the production of the charmonium is of great interest for the theory due to its richness of $c\bar{c}$ states, the properties of which have not been sufficiently studied. We know that a lot of charmonia have been observed since 1974 A18 ; A19 , but this is not the end of the whole history of the charmonium family, especially of observations of charmoniumlike XYZ states A20 ; A21 .
It is clear that the charmonium family group, by adding more states, may provide important information for understanding the nonperturbative behavior of the strong interaction, which is intensively associated with color restriction in quantum chromodynamics (QCD). The vector charmoniumlike $Y$ state plays a special role in the construction of the charmonium family.
@@
We also would like to note that in experimental measurements, all charmonium states were observed below the threshold of a charmed pair of quarks in mass ($2m_{D}$), and it was shown that their observed spectrum is consistent with the prediction of the model of the potential of charm and anticharm A27 .
Two-particle baryon decays of vector ($J^{PC}=1^{--}$) charmonium (like) resonances provide a rich laboratory for predictions of quantum chromodynamics A28 ; A29 .
@@
For the process $e^{+}e^{-}\to\bar{p}p$, the authors of the work Bianconi noted that there are deviations in the existence in the $BABAR$ data in timelike proton form factors from the pointlike behavior of proton-antiproton electromagnetic current. Namely, in this paper $F_{0}$ + $F_{osc}$ was used in terms of the form factor where $F_{0}$ is expressed as the long-term trend of the form factor, and $F_{osc}$ is a function that has the form exp(-Bp)cos(Cp) where $p$ is the relative momentum of the $\bar{p}p$ pair in the final state.
@@
In these experimental works A8 ; I5 ; I6 ; I8 ; I9 ; I10 ; I11 ; I13 , from the point of view of available data, there is an electromagnetic form factor in a timelike region, which is consistent with naive rules for counting quarks and the prediction of perturbative QCD (pQCD) at large $q^{2}$ I14 ; I15 . Therefore, in these cases, to provide a unique opportunity to study the internal structure of hadrons and the electromagnetic properties of hyperons, timelike form factors are used.
It should be noted that the data listed in I16 , which are obtained mainly from experiments on electron-positron and proton-proton (proton-antiproton) interactions, determine today most of our knowledge about experimentally established nucleon resonances. The charmonium resonance mass is present in the transition region between the perturbative and nonperturbative regimes of the $\psi(3770)$. To gain knowledge about the structure in perturbative and nonperturbative strong interactions, it is necessary to study the strong and hadronic decays of $\psi(3770)$ in this energy region I17 .
After analyzing these samples together with the data, it became possible to conduct a theoretical study taking into account the interference of resonant and nonresonant amplitudes for the exclusive decay of $\psi(3770)$. It is known that for several years BESIII has collected the largest sample of data on $e^{+}e^{-}$ collisions at an energy of 3.773 GeV.
@@
What follows is the decay process in these charmonium states, i.e., the decay into light hadrons via either a three-gluon process ($e^{+}e^{-}\to\psi\to ggg\to hadrons$) or a one-photon process ($e^{+}e^{-}\to\psi\to\gamma^{*}\to hadrons$).
Really, according to the Okubo-Zweig-Iizuki rule (OZI), it is expected that $\psi(3770)$, the lowest charmonium state lying at $1^{--}$ above the threshold of $D\bar{D}$, will predominantly decay into the $D\bar{D}$ final states A17 ; I18 ; I20 . To test a QCD prediction that can be understood in terms of the amplitudes of the quark distribution in hadron-hadron pairs and the conservation of the total hadron helicity, it is necessary to study the process of $\psi(3770)$ production in $e^{+}e^{-}$ annihilation and its subsequent decay into two hadrons.
@@
It can be expected that the resonance of $\psi(3770)$ will almost completely decay into pure $D\bar{D}$ A17 , and in the baryon sector at the electron-positron collider, the production of a baryon-antibaryon pair can be tested by fundamental symmetries, in particular, when the probability of the process increases due to the resonance such as the $J/\psi$ (or $\psi$)I22 .
It should be noted that the BESIII collaboration has until now conducted highly accurate studies of the possible threshold increase in the $e^{+}e^{-}\to\Sigma^{\pm}\bar{\Sigma}^{\mp}$ I23 , $\Xi^{-}\bar{\Xi}^{+}$ I24 ; I24A and $\Sigma^{0}\bar{\Sigma}^{0}$ I25 processes and has also shown that the cross section does not vanish near the threshold.
It can be understood that the threshold effect thus obtained will be useful for measuring the near-threshold pair generation of hyperons in the production of $\Sigma^{+}\bar{\Sigma}^{-}$ in the $e^{+}e^{-}$ annihilation process, which was observed in the BESIII experiment I26 .
@@
Our aim in this paper is to study the characteristics of the $\Sigma^{+}\bar{\Sigma}^{-}$ production in the $e^{+}e^{-}$ annihilation process, including beyond the one-photon exchange also the exchange of several charmonium(like) states, $\psi(3770),\,\psi(4040),\,\psi(4160),\,Y(4230),\,Y(4360),\,\psi(4415),\,\psi(4660)$.
@@ II The process $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ in Born approximation
$d\Phi_{2}=(2\pi)^{4}\delta(p_{+}+p_{-}-q_{+}-q_{-})\frac{d\bf{q_{+}}}{(2\pi)^{3}2E_{+}}\frac{d\bf{q_{-}}}{(2\pi)^{3}2E_{-}}=$
$=\frac{1}{(2\pi)^{2}}\delta(p_{+}+p_{-}-q_{+}-q_{-})\frac{d\bf{q_{+}}}{2E_{+}}\frac{d\bf{q_{-}}}{2E_{-}}=\frac{|\bf{p}|}{16\pi^{2}\sqrt{s}}\,d\Omega_{\Sigma}=\frac{\beta}{16\pi}\,d\cos\theta_{\Sigma},$ (5)
where $d\Omega_{\Sigma}=d\phi_{\Sigma}\,d\cos\theta_{\Sigma}=2\pi\,d\cos\theta_{\Sigma}$, and $\phi_{\Sigma}$ and $\theta_{\Sigma}$ are the azimuthal and polar angles of the $\Sigma$-hyperon momentum in the $e^{+}e^{-}$ rest frame, that is $\theta_{\Sigma}$ is the angle between the directions of the momenta of the electron $\bf{p_{-}}$ and $\Sigma$-hyperon $\bf{q}_{+}$ (Fig. 1). We note that the modulus of three momenta of the final $\Sigma^{+}$-hyperon (or $\bar{\Sigma}^{-}$-hyperon) is defined in this case by the $\delta$-function in the phase volume, i.e.,
$|{\bf{p}}|\equiv|{\bf{q_{+}}}|=|{\bf{q_{-}}}|=\frac{\sqrt{s}}{2}\sqrt{1-\frac{4M^{2}_{\Sigma}}{s}}=\frac{\sqrt{s}}{2}\beta,$ (6)
@@
where $C$ is a free parameter that fits the experimental data and $\Lambda_{QCD}$ is the QCD scale parameter. We want to note that the constant $C$ should be fitted for the hyperon-antihyperon production according to specific experimental data in the energy range of the corresponding experiment.
In the present work for the $\Sigma^{+}\bar{\Sigma}^{-}$ pair production in the $e^{+}e^{-}$ annihilation process (1), we fix this constant $C$ using the BESIII measurement I26 , and in Fig. 3 we show the dependence of the total cross section as a function of the center-of-mass energy $\sqrt{s}$. When the value is $\Lambda_{QCD}$ = 300 MeV in the Born cross section from (11) for the parameter $C$, after fitting with respect to these data we obtain the following value:
@@ III The quarkonium $\psi(3770)$ intermediate state
but only with a different constant $g_{e}=F_{1}^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}(M_{\psi}^{2})$, that depends on the charmonium mass and is equal to the value of the form factor $F_{1}^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}(M_{\psi}^{2})$ which is the value of the form factor for the vertex $\psi\to\Sigma^{+}\bar{\Sigma}^{-}$ on the $\psi(3770)$ mass shell [we note that here we accept the same approximation as in the Born case and assume that $F_{2}^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}(M_{\psi}^{2})=0]$. We know from I16 the total decay width of $\psi\to e^{+}e^{-}$, which is equal to $\Gamma_{\psi\to e^{+}e^{-}}=261\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{eV}$, and we can calculate this constant $g_{e}$,
@@
It is shown Kuraev that the imaginary part of the vertex $e\bar{e}\to\psi$ is small and it is less than 10 % of the real part. In this case, we neglect such a possible imaginary part and we do not include this source of error when computing the statistical uncertainty of our final results.
@@
$\approx-eg_{e}\,\mbox{Tr}[\hat{p_{-}}\gamma_{\mu}\hat{p_{+}}\gamma^{\mu}]\approx 4\,e\,g_{e}s.$ (23)
@@
In this formula (24), the subintegral expression contains the dynamics of the transformation of the charmonium into the $\Sigma^{+}\bar{\Sigma}^{-}$ pair, and therefore, it can be written in a separate form
$S_{i}(s)=\frac{eg_{e}\beta}{48\pi s}\int\limits_{-1}^{1}d\cos\theta_{\Sigma^{+}}\sum_{s^{\prime}}(J_{\gamma\to\Sigma^{+}\bar{\Sigma}^{-}}^{\alpha})^{*}J^{\psi\to\Sigma^{+}\bar{\Sigma}^{-}}_{\alpha}.$ (25)
Thus, if we substitute formula (25) into formula (24), then the expression for the interference contribution to the total cross section can be written in the following form:
@@
In order to reconstruct the total cross section, we use the procedure described in Bystritskiy [Eqs. (15) and (16)] and with the help of the interference contribution (26) with the total relative phase between the Born amplitude $\mathcal{M}_{B}$ and the amplitude with the charmonium contribution $\mathcal{M}_{\psi}$.